Two points (p,q) on a linear supply equation are (3.5,116000) and (5,180000)

Determine supply function.


1
Expert's answer
2021-03-09T03:57:55-0500

compose the equation of a line at two points:

"\frac{{p - 3.5}}{{5 - 3.5}} = \frac{{q - 116000}}{{180000 - 116000}} \Rightarrow \frac{{p - 3.5}}{{1.5}} = \frac{{q - 116000}}{{{\rm{64000}}}} \Rightarrow q - 116000 = \frac{{{\rm{64000}}}}{{1.5}}\left( {p - 3.5} \right) \Rightarrow"

"\Rightarrow q = \frac{{128000}}{3}p - \frac{{100000}}{3} \Rightarrow q = \frac{{128000p - 100000}}{3}"


Answer: "q = \frac{{128000p - 100000}}{3}"